Circular No. 8988

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2008 T1 (BOATTINI) A. Boattini reports his discovery of another comet on CCD images

taken with the Mt. Lemmon 1.5-m reflector (discovery observation tabulated below), the fairly condensed coma being ∼ 10′′ in diameter, with a fan tail ∼ 60′′ long spanning p.a. 245^{o}–275^{o} (with the brightest tail structure in p.a. 250^{o}). Four stacked 60-s follow-up unfiltered images with the same telescope by R. E. Hill on Oct. 2.4 UT show a coma diameter of 10′′-12′′ and a broad, diffuse tail of length 40′′–50′′ in p.a. ∼ 300^{o}. Following posting on the ‘NEOCP’ webpage, P. Birtwhistle (Great Shefford, Berkshire, England, 0.40-m reflector, Oct. 2.1) finds a diffuse coma of diameter 10′′, extending 15′′ in p.a. 248^{o} (with a possible straight tail extending for 150′′ at the same p.a.). R. Ligustri (Talmassons, Udine, Italy, 0.25-m reflector, Oct. 2.2) reports that seven 300-s co-added exposures show a coma diameter of ∼ 15′′. J. E. McGaha (Tucson, AZ, U.S.A., 0.62-m reflector, Oct. 2.2) writes that a 240-s exposure shows a bright nuclear condensation of diameter 5′′ with a faint coma of diameter 12′′ and a faint tail 150′′ long in p.a. 256^{o}. Images taken by R. Holmes (Charleston, IL, U.S.A., 0.61-m astrograph, Oct. 2.3) and measured by H. Devore show a wide tail 2′01′′ long in p.a. 251^{o}.6. J. Young (Table Mountain 0.61-m reflector) finds a very small bright coma of diameter 6′′ with a hint of a broad, weak, fan-shaped tail ∼ 16′′ long spanning 225^{o}–310^{o}. E. Guido, Castellammare di Stabia, Italy, writes that eighty stacked 60-s unfiltered CCD exposures taken remotely by G. Sostero, P. Camilleri, and himself with a 0.25-m reflector near Mayhill, NM, U.S.A., on Oct. 2.4 show a coma ∼ 10′′ in diameter and a broad, fan-shaped tail ∼ 40′′ toward p.a. 244^{o}.

2008 UT  α_{2000}  δ_{2000}  Mag.
Oct. 1.34535  0^{h}42^{m}43.^{s}52  +7^{o}48′45′′.6  18.0

The available astrometry (including prediscovery Spacewatch astrometry from Sept. 2 and 21), the following elliptical orbital elements by B. G. Marsden, and an ephemeris appear on MPEC 2008-T30. Marsden notes that the comet made a close approach to Jupiter in Apr. 2003 (minimum distance 0.03 AU).

T = 2008 Feb. 25.3012 TT  ω = 35.^{o}4139
e = 0.287660  Ω = 291.6541 }2000.0
q = 3.028416 AU  i = 2.0806
a = 4.251365 AU n^{o} = 0.1124376 P = 8.77 years
2008 October 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green